#1problems: a titration is performed by adding 0.600M KOH to 40.0 ML of 0.800 M HCL.. calculate the ph before addition of any koh.. calculate the ph after addition of 5.0 ML of the base..

 
#1problems:
a titration is performed by adding 0.600M KOH to 40.0 ML of 0.800 M HCL..
  1. calculate the ph before addition of any koh..
  2. calculate the ph after addition of 5.0 ML of the base..

Grade:12th pass

1 Answers

Sunil Kumar FP
9 years ago
equation
koh+hcl----kcl + h2o
millimole of hcl=40*.8=32
millimole of koh=.6*5=3
ph before addition of koh=-log(h+)=-log.8=.097
after addition of koh=hcl remains=32-3=29miilimole
volume of the solution=45ml
concentration of hcl=29/45=.644
log(.644)=.191
thanks and regards
sunil kr
askIITian faculty